# Amorphous Solid Model of Vectorial Hopfield Neural Networks

### v4

@@ Abstract
We introduce a three-dimensional vectorial extension of the Hopfield associative-memory model in which each neuron is a unit vector on $S^{2}$ and synaptic couplings are $3\times 3$ blocks generated through a vectorial Hebbian rule. The resulting block-structured operator is mathematically analogous to the Hessian of amorphous solids and induces a rigid energy landscape with deep minima for stored patterns. Simulations and spectral analysis show that the vectorial network substantially outperforms the classical binary Hopfield model. For moderate connectivity, the critical storage ratio $\gamma_{c}$ grows approximately linearly with the coordination number $Z$, while for $Z\gtrsim 40$ a high-connectivity regime emerges in which $\gamma_{c}$ systematically exceeds the extrapolated low-$Z$ linear fit. At the same time, a persistent spectral gap separates pattern modes from the bulk and basins of attraction enlarge, yielding enhanced robustness to initialization noise. Thus geometric constraints combined with amorphous-solid-inspired structure produce associative memories with superior storage and retrieval performance, especially in the high-connectivity ($Z\gtrsim 20$$30$) regime.
@@
The Hopfield model Hopfield (1982) is a cornerstone of associative memory in neural networks and artificial intelligence Meshulam and Bialek (2025). In its classical formulation, NN binary neurons $s_{i}\in\{-1,+1\}$ evolve according to symmetric interactions $W_{ij}$ defined by the Hebbian rule from stored patterns. The dynamics minimizes $E=-\frac{1}{2}\sum_{ij}W_{ij}s_{i}s_{j}$, leading to pattern retrieval via convergence to energy minima.
Connections between neural network models and disordered materials have been explored in spin glasses Amit et al. (1985); Nicoletti et al. (2025), jammed packings Geiger et al. (2019); Zhang et al. (2025), random resistor and memristive networks Hu et al. (2015), and elastic networks Altman et al. (2024). Recent work has emphasized the relation between the “cost landscape” and the physical landscape encoded by the cost Hessian and physical Hessian, respectively Stern et al. (2025).
@@
The vectorial extension is relevant to physical systems where orientational degrees of freedom are essential, such as liquid crystals, polymeric glasses, and biological networks (e.g. the cytoskeleton). It reformulates the Hopfield framework in terms of block-structured random matrices, enabling systematic comparison of learning efficiency and computational properties with classical binary models Hertz et al. (1991). We start by reviewing the binary Hopfield model.
We consider 3D networks of NN nodes (neurons) connected by 3D vectors (“bonds” or synapses) as in Fig. 1. In the amorphous solid analogy, the connections are springs with spring constant $\kappa$ Milkus and Zaccone (2016, 2017); Zaccone (2023).
Figure 1: Rendering of a 3D network of NN nodes with average coordination number $Z=7$ and nearly uniform orientation distribution.
The binary Hopfield model consists of NN neurons with $s_{i}\in\{-1,+1\}$. To store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ with $\boldsymbol{\xi}^{\mu}=(\xi_{1}^{\mu},\ldots,\xi_{N}^{\mu})$ and $\xi_{i}^{\mu}\in\{-1,+1\}$, the weight matrix is
@@
In the vectorial model, each neuron is a unit vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ with $|\mathbf{s}_{i}|=1$. We store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$, where each pattern is a set of NN random unit vectors on $S^{2}$:
@@
As a first test, we verified that the vectorial network correctly retrieves stored memories. From random initial configurations, the dynamics converges to memorized attractors (or complementary states) in fewer than $\sim 30$ asynchronous updates for all trials, showing that stored configurations are stable fixed points of the energy landscape. This is, to our knowledge, the first working implementation of a three-dimensional vectorial Hopfield model.
@@
which satisfies $W\Xi=\Xi$, so each stored pattern is an exact eigenvector of $W$ with eigenvalue 11. This guarantees perfect recall of all training patterns in the fully connected case and provides a natural extension of the classical formulation to correlated patterns.
@@
where $X_{i}$ collects neighbour states and $Y_{i}$ the target states for neuron ii. The parameter $\lambda$ ensures numerical stability and controls smoothness. This local pseudo-inverse generalizes the fully connected rule to sparse graphs while encoding topology.
@@
(b) Spectral estimate
Figure 2: Critical storage capacity versus connectivity. (a) Dynamical estimate of $\gamma_{c}$ with standard-deviation error bars across 55 replicas. (b) Spectral estimate from the vanishing smallest eigenvalue of the Hessian at each attractor. Both panels show an approximately linear dependence of $\gamma_{c}$ on $Z$ with closely matching slopes.
@@
The vectorial Hopfield model generates energy landscapes with multiple local minima corresponding to stored patterns. We quantify pattern stability via the energy difference between stored and random configurations. For $N=25$ and $\gamma=0.1$, stored patterns have mean energies about 77 units lower than random states, ensuring robust retrieval against initialization noise. The energy function
naturally creates this hierarchy through the outer-product construction of $\mathbf{W}$. The energy gap decreases as $\gamma$ approaches $\gamma_{c}$, paralleling metastable hierarchies in glass-forming systems Biroli and Kurchan (2001); Han et al. (2020); Charbonneau et al. (2014).
The eigenvalue spectrum of $W_{ij}$ [Eq. (1)] characterizes computational properties at different pattern densities (see Fig. 5 in the Appendix). For low density ($\gamma=0.05$), the spectrum shows isolated pattern eigenvalues and a bulk near zero. For intermediate density ($\gamma=0.1$$0.2$), a bulk develops with pattern eigenvalues as outliers. For high density ($\gamma=0.5$), the spectrum approaches random-matrix behavior. The evolution of spectral density with $\gamma$ reflects a crossover from a memory-dominated to a noise-dominated regime; $\gamma_{c}$ corresponds to the merging of pattern eigenvalues with the bulk and the loss of discriminability.
Convergence properties depend on initialization and pattern density. The convergence time scales logarithmically with system size NN for successful retrievals, while basin size decreases approximately exponentially as $\gamma$ approaches $\gamma_{c}$, and basin boundaries become more complex. The unit-vector constraint produces dynamics on the product manifold $S^{2}\times S^{2}\times\cdots\times S^{2}$, preserving pattern structure but adding geometric complexity. The update rule
@@
More broadly, our findings highlight the usefulness of importing concepts from elasticity and glassy physics into neural-network design. Promising extensions include combining vectorial representations with nonlinear learning rules, exploring sparse or hierarchical connectivities that retain the high-$Z$ advantage, and generalizing to higher-dimensional state manifolds Krotov and Hopfield (2021). Overall, the vectorial Hopfield framework provides a compact, high-capacity, and dynamically robust associative memory that markedly outperforms its classical binary counterpart, with particularly strong gains in the the moderate to high-connectivity regime.
@@ Acknowledgments
AZ gratefully acknowledges funding from the US Army DEVCOM Army Research Office through contract nr. W911NF-22-2-0256 and from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. During the preparation of this work we became aware of similar observations in a vectorial spin-glass model Nicoletti et al. (2025).
appendix
In Fig. 5, we show the eigenvalue distribution of $W_{ij}$ for the 3D networks proposed here, obtained by direct numerical diagonalization of Eq. (1). In Fig. 5 we show the typical time evolution of the Lyapunov energy as a function of the number of update steps, illustrating the monotonic convergence of the dynamics.
Figure 5: Evolution of the spectral density $\rho(\lambda)$ of the weight matrix $W_{ij}$ with pattern capacity $\gamma=P/(Nd)$. Low density shows a concentrated band near zero with isolated positive eigenvalues, while high density exhibits broad continuous distributions indicating pattern interference. The spectra are obtained by direct numerical diagonalization of $W_{ij}$ as defined in Eq. (1) for the 3D networks of Fig. 1.
0551010151520208-86-64-42-20Update stepEnergy
Figure 6: Time evolution of the Lyapunov energy $E(t)=-\tfrac{1}{2}\,\mathbf{s}^{\!\top}\!W\,\mathbf{s}$ for a single synchronous run of the vector–Hopfield model with $N=25$ neurons, $d=3$, and $P=5$ stored patterns. After a rapid decay in the first few updates the energy settles on the minimum corresponding to the retrieved pattern, confirming monotonic convergence of the dynamics.

### v5

@@ Abstract
We introduce a three-dimensional vector Hopfield model where each neuron is a unit vector on $S^{2}$ and synaptic weights are $3\times 3$ blocks built by a vectorial Hebbian rule, analogous to the Hessian of amorphous solids. The resulting rigid energy landscape strongly stabilizes stored patterns. Simulations and spectral analysis show that the critical storage ratio grows roughly linearly with connectivity and enters a high-connectivity regime with enhanced capacity, a persistent spectral gap, and enlarged basins of attraction, yielding robust associative memory.
@@
The Hopfield model Hopfield (1982) is a cornerstone of associative memory in neural networks and artificial intelligence Meshulam and Bialek (2025). In its classical formulation, $N$ binary neurons $s_{i}\in\{-1,+1\}$ evolve according to symmetric interactions $W_{ij}$ defined by the Hebbian rule from stored patterns. The dynamics minimizes $E=-\frac{1}{2}\sum_{ij}W_{ij}s_{i}s_{j}$, leading to pattern retrieval via convergence to energy minima. Connections between neural network models and disordered materials have been explored in spin glasses Amit et al. (1985); Baity-Jesi et al. (2018); Nicoletti et al. (2025), jammed packings Geiger et al. (2019); Zhang et al. (2025), random resistor and memristive networks Hu et al. (2015), and elastic networks Altman et al. (2024). Recent work has emphasized the relation between the “cost landscape” and the physical landscape encoded by the cost Hessian and physical Hessian, respectively Stern et al. (2025).
@@
The vectorial extension is relevant to systems with orientational degrees of freedom, such as liquid crystals, polymeric glasses, biological networks such as the cytoskeleton, and enables systematic comparison with classical binary Hopfield networks Hertz et al. (1991). We start by reviewing the binary Hopfield model.
We consider 3D networks of $N$ nodes (neurons) connected by 3D vectors (“bonds” or synapses) as in Fig. 1. In the amorphous solid analogy, the connections are springs with spring constant $\kappa$ Milkus and Zaccone (2016, 2017); Zaccone (2023).
Figure 1: Rendering of a 3D network of $N$ nodes with average coordination number $Z=7$ and nearly uniform orientation distribution.
The binary Hopfield model consists of $N$ neurons with $s_{i}\in\{-1,+1\}$. To store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ with $\boldsymbol{\xi}^{\mu}=(\xi_{1}^{\mu},\ldots,\xi_{N}^{\mu})$ and $\xi_{i}^{\mu}\in\{-1,+1\}$, the weight matrix is
@@
In the vectorial model, each neuron is a unit vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ with $|\mathbf{s}_{i}|=1$. We store $P$ patterns $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$, where each pattern is a set of $N$ random unit vectors on $S^{2}$:
@@
Simulations show reliable retrieval: from random initial configurations, dynamics converges to stored patterns (or complements) within $\sim 30$ asynchronous updates.
@@
which satisfies $W\Xi=\Xi$, so each stored pattern is an exact eigenvector of $W$ with eigenvalue $1$. This guarantees perfect recall of all training patterns in the fully connected case and provides a natural extension of the classical formulation to correlated patterns.
@@
where $X_{i}$ collects neighbour states and $Y_{i}$ the target states for neuron $i$. The parameter $\lambda$ ensures numerical stability and controls smoothness. This local pseudo-inverse generalizes the fully connected rule to sparse graphs while encoding topology.
@@
Figure 2: Critical storage capacity versus connectivity. Top figure: Dynamical estimate of $\gamma_{c}$ with standard-deviation error bars across 55 replicas. Bottom figure: Spectral estimate from the vanishing smallest eigenvalue of the Hessian at each attractor. Both panels show an approximately linear dependence of $\gamma_{c}$ on $Z$ with closely matching slopes.
@@
The vectorial Hopfield model generates energy landscapes with multiple local minima corresponding to stored patterns. We quantify pattern stability via the energy difference between stored and random configurations. For $N=25$ and $\gamma=0.1$, stored patterns have mean energies about $7$ units lower than random states, ensuring robust retrieval against initialization noise. The energy function given by Eq. (6) naturally creates this hierarchy through the outer-product construction of $\mathbf{W}$. The energy gap decreases as $\gamma$ approaches $\gamma_{c}$, paralleling metastable hierarchies in glass-forming systems Biroli and Kurchan (2001); Han et al. (2020); Charbonneau et al. (2014).
The eigenvalue spectrum of $W_{ij}$ [Eq. (1)] characterizes computational properties at different pattern densities (see Fig. 7 in the End Matter). For low density ($\gamma=0.05$), the spectrum shows isolated pattern eigenvalues and a bulk near zero. For intermediate density ($\gamma=0.1$$0.2$), a bulk develops with pattern eigenvalues as outliers. For high density ($\gamma=0.5$), the spectrum approaches random-matrix behavior. The evolution of spectral density with $\gamma$ reflects a crossover from a memory-dominated to a noise-dominated regime; $\gamma_{c}$ corresponds to the merging of pattern eigenvalues with the bulk and the loss of discriminability.
Convergence properties depend on initialization and pattern density. The convergence time scales logarithmically with system size $N$ for successful retrievals, while basin size decreases approximately exponentially as $\gamma$ approaches $\gamma_{c}$, and basin boundaries become more complex. The unit-vector constraint produces dynamics on the product manifold $S^{2}\times S^{2}\times\cdots\times S^{2}$, preserving pattern structure but adding geometric complexity. The update rule
@@
Promising extensions include nonlinear learning rules, sparse or hierarchical connectivities that retain the high-$Z$ advantage, and higher-dimensional state manifolds Krotov and Hopfield (2021). Overall, the vectorial Hopfield framework provides a compact, high-capacity associative memory that markedly outperforms the classical binary model in the moderate- to high-connectivity regime.
@@ Acknowledgments
AZ gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. During the preparation of this work we became aware of similar observations in a vectorial spin-glass model Nicoletti et al. (2025).
End Matter
This appendix collects the numerical diagnostics supporting the results presented in the main text. We report: (i) direct evidence of Lyapunov convergence; (ii) the spectral softening of the Hessian leading to the critical load; (iii) the evolution of the weight–matrix spectrum with increasing pattern density; (iv) quantitative energy separation between memory and spurious minima.
01010202030-3020-2010-100Update stepEnergy
Figure 5: Monotonic decay of the Lyapunov energy $E(t)$ during synchronous updates for $N=25$, $d=3$, $P=5$. The system converges rapidly to the minimum associated with the retrieved memory.
The loss of stability of stored patterns is captured by the smallest eigenvalue $\lambda_{\min}$ of the Hessian. The curve in Fig. 6 shows the softening and zero-crossing of $\lambda_{\min}$ as the load $\gamma$ approaches the critical value.
[…]
Figure 6: Minimum eigenvalue of the Hessian for a fully connected network ($Z=63$). The zero-crossing at $\gamma_{c}^{\text{spec}}\simeq 0.21$ marks the onset of instability of the memory attractors.
Figure 7 displays the spectral density of the weight matrix $W$ for increasing pattern loads. As $\gamma$ grows, isolated outliers merge with the bulk and the spectrum broadens into a continuous band, consistent with the breakdown of retrieval capacity.
Figure 7: Evolution of the spectral density $\rho(\lambda)$ of the weight matrix with increasing pattern load $\gamma$. Low loads (top left) exhibit isolated pattern eigenvalues; intermediate loads develop a broadened bulk; high loads (bottom right) show a continuous spectrum approaching random-matrix form.
Finally, the table below summarises the energy separation between stored patterns and spurious minima for three representative load levels. Energies are obtained from 300 random initialisations for each value of $\gamma/\gamma_{c}$.
$\gamma/\gamma_{c}$ $E_{\min}^{\text{mem}}$ $E_{\min}^{\text{sp}}$ Gap $\Delta E$
0.1 30.397-30.397
0.3 27.681-27.681 27.199-27.199 +0.482+0.482
0.7 24.952-24.952 24.884-24.884 +0.068+0.068
Table 1: Energies of stored memories and deepest spurious minima. Large gaps at low load reflect clean memory dominance; near capacity the gap collapses, consistent with interference-driven instability.